Problems and Applications Q3

Frances and Latasha are roommates. They spend most of their time studying (of course), but they leave some time for their favorite activities: making pizza and brewing root beer. Frances takes 5 hours to brew a gallon of root beer and 3 hours to make a pizza. Latasha takes 8 hours to brew a gallon of root beer and 4 hours to make a pizza.

Frances's opportunity cost of brewing a gallon of root beer is   , and Latasha's opportunity cost of brewing a gallon of root beer is   .

  has an absolute advantage in brewing root beer, and   has a comparative advantage in brewing root beer.

If Frances and Latasha trade foods with each other,   will trade away pizza in exchange for root beer.

The price of pizza can be expressed in terms of gallons of root beer. The highest price at which pizza can be traded that would make both roommates better off is   of root beer, and the lowest price that makes both roommates better off is   of root beer per pizza.
Answer #1

France:

Opportunity cost of brewing a gallon of beer = 5/3 = 1.67 pizza

Latasha:

Opportunity cost of brewing a gallon of beer = 8/4 = 2 pizza

---

Comparative advantage in beer = One who has lower opportunity cost of brewing beer = France

Absolute advantage in beer = France

Reason: Suppose there are 120 hours available for work.

France will produce = 120/5 = 24 beer while Latasha will produce = 120/8 = 15 beer

----

Latasha will trade away pizza for root beer.

Reason: Since Latasha has comparative advantage in pizza production, she will produce pizza and sell it in exchage for beer

----

Highest price: 2 pizza

Lowest price: 1.67 pizza
Answer #2
Frances and Latasha are roommates. They spend most of their time studying (of course), but they leave some time for their favorite activities: making pizza and brewing root beer. Frances takes 3 hours to brew a gallon of root beer and 2 hours to make a pizza. Latasha takes 7 hours to brew a gallon of root beer and 5 hours to make a pizza. Frances's opportunity cost of making a pizza is 2/3 gallon    of root beer, and Latasha's opportunity cost of making a pizza is5/7 gallon of root beer. Frances has an absolute advantage in making pizza, and Frances has a comparative advantage in making pizza. The price of pizza can be expressed in terms of gallons of root beer. The highest price at which pizza can be traded that would make both roommates better off is 5/7 gallon  of root beer, and the lowest price that makes both roommates better off is 2/3 gallon of root beer per pizza. If Frances and Latasha trade foods with each other, Frances will trade away pizza in exchange for root beer.
source: Cengage
answered by: Purple Hyacinth01